OpenStudy (anonymous):

. Find the product: (7x²-8y)(7x²+8y)

OpenStudy (anonymous):

i got 49x⁴-56xy+56xy+64y² 49x⁴-112xy+64y²

hero (hero):

Hint: Difference of Squares

OpenStudy (anonymous):

Check that middle term: "-56xy + 56xy"

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Also (-8y)*(+8y) = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Try again. It is multiplication, not addition.

OpenStudy (anonymous):

ok im lost now

OpenStudy (anonymous):

Start over.

OpenStudy (anonymous):

\[\large (7x²-8y)(7x²+8y)=(7x^2)(7x^2)+(7x^2)(8y)+(-8y)(7x^2)+(-8y)(8y)\]

OpenStudy (anonymous):

multiply the first terms of each expression together...write down then multipy the first term of the expression by the second term of the 2d exp...write down then multiply the 2d term of the first exp by the first term of the 2d expression..write it down finally multipy th 2s term of the 1st exp by the 2d term in the 2s expression...write it down simplify combining like terms and you have the answer. or FOIL -----> First terms, Outside terms, Inside terms, Last terms add them together after multiplying

OpenStudy (anonymous):

\[\left(7 x^2-8 y\right) \left(7 x^2+8 y\right)=49 x^4-64 y^2 \]